OpenStudy (anonymous):

Plsee answer A and B enters a line of 20 people randomly. What is the probability that there are 6 people standing between A and B?

OpenStudy (kropot72):

Drawing a diagram shows that there are 30 permutations where A and B have 6 people between them. The number of permutations of A and B in the total of 22 people is \[22P2=\frac{22!}{20!}\] therefore the required probability is given by \[30\times \frac{20!}{22!}=you\ can\ calculate\]